OpenStudy (anonymous):

A car is traveling on a highway. The distance (in miles) from its destination and the time (in hours) is given by the equation d=420-65t. What is the practical meaning of the d-intercept? a. Initially the car is 355 miles from its destination. b. Initially the car is 420 miles from its destination. c. Initially the car is 65 miles from its destination. d.Initially the car is 485 miles from its destination.

OpenStudy (anonymous):

Which number is a variable and which number is a constant in your equation?

OpenStudy (anonymous):

Variable?

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

65t is a variable as it will change as t changes.

OpenStudy (anonymous):

So D is the constant?

OpenStudy (anonymous):

No d is the solution and will change as t changes.

OpenStudy (anonymous):

Okay, now what? (Im sorry, thank you so much)

OpenStudy (anonymous):

Let t represent time in hours and D represent distance traveled. So of we are starting a trip the first value for t would be 0, because no time has passed. Thus when t = 0 this will represent where we are starting our trip from. distance wise.

OpenStudy (anonymous):

d=420-65(0) ?

OpenStudy (anonymous):

So d= what at our starting point?

OpenStudy (anonymous):

I have no clue! :(

OpenStudy (anonymous):

I don't completely understand what you are asking

OpenStudy (calculusfunctions):

An equation of a line is of the form y = mx + b, where m is the slope (the rate of change of y with respect to x) and b is the y-intercept (the initial value). You can now compare your equation to this one to figure it out. Hint: in your equation d is the same as y and t is the same as x.

OpenStudy (anonymous):

d=420-65t y=mx+b I still am confused

OpenStudy (calculusfunctions):

d = 420 - 65t y = b + mx If b is the initial value, then what is b in your equation? This is the initial distance from the destination.

OpenStudy (anonymous):

420?

OpenStudy (calculusfunctions):

You can also look at it another way. Initial distance d is when time t = 0. Thus you can substitute t = 0 in the equation and solve for d to determine the initial distance.

OpenStudy (calculusfunctions):

Yes! You got it! The initial distance is 420.

OpenStudy (anonymous):

Lol, it B? If it is i figured that out right before you sent it

OpenStudy (calculusfunctions):

Yes it's B

OpenStudy (anonymous):

Thank you!!!!!!!!!!

OpenStudy (calculusfunctions):

You're welcome.